# Submitted toTopology Proceedings

### v2

@@ 5. Heteroclinic regions
In this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, and at least two equilibria $a_{+},a_{-}\in\mathbb{C}{}$ of (1.1) with at least one heteroclinic orbit connecting these two equilibria. In particular, we assume that $a_{\pm}$ is not a center. Under these assumptions, it is possible to investigate topologically another characteristic set, the so-called "heteroclinic region". Taking up the questions at the end of chapter 3, a heteroclinic region is one possible structure between two adjacent global elliptic sectors. We analyse it topologically and illustrate it by an example.
@@
+():=$\mathcal{H}_{+(-)}:=\left\{x\in\mathbb{C}{}:w_{+(-)}(\Gamma(x))=\{a_{+(-)}\}\right\}\text{.}$
@@
Both, $a_{+}$ as well $a_{-}$ is either a focus or node or an equilibrium of order $m\in\mathbb{N}\setminus\{1\}$. Let $x\in A_{+}$ and assume that $a_{+}$ is an equilibrium of order $m\geq 2$. By [10, Proposition 4.3], there exist numbers $r,\theta,\Theta,\tau>0$ such that for all $z\in\{\hat{z}\in\mathbb{C}{}:|\hat{z}-a_{+}|<r,|\arg(\hat{z}-a_{+})-\theta|<\Theta\}=:A$ the orbit through $z$ tends to $a$ for $t\to\infty$ and $M:=\Phi(\tau,x)\in A$. Since $A$ is open, we can choose $\varepsilon>0$ small enough such that $\mathcal{B}_{\varepsilon}(M)\subset A$. By continuity of the flow, cf. [19, Chapter 2.4, Theorem 4], there exists $\delta>0$ such that $|\Phi(\tau,x)-\Phi(\tau,y)|<\varepsilon$ for all $y\in\mathcal{B}_{\delta}(x)$, i.e. $\mathcal{B}_{\delta}(x)\subset\mathcal{H}_{+}$ and $\mathcal{H}_{+}$ is open. Additionally, if $a_{+}$ is a node or focus, then we can apply Theorem 4.3 to conclude again the openness of $\mathcal{H}_{+}$. Since the same argumentation holds for $a_{-}$, $\mathcal{H}_{-}$ is also open. It follows that $\mathcal{H}=\mathcal{H}_{+}\cap\mathcal{H}_{-}$ is open. ∎
Lemma 5.4.
The basins of all centers, foci and nodes $a\in F^{-1}(\{0\})\setminus\{a_{+},a_{-}\}$ have empty intersection with $\mathcal{H}$. In particular, there are no centers, nodes and foci on $\partial\mathcal{H}$.
Proof.
There are no periodic orbits in $\mathcal{H}$. Additionally, $a\not\in\{a_{+},a_{-}\}$. Thus, this proof is analogous to the proof of Lemma 4.4. ∎
@@
Since there exists at least one heteroclinic orbit in $\mathcal{H}$ connecting $a_{+}$ and $a_{-}$, we have $\{a_{+},a_{-}\}\subset\partial\mathcal{H}$. Suppose there exists an equilibrium $a\in\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. By Lemma 5.4, $a$ is an equilibrium of order $m\geq 2$, i.e. $a$ has a FED, cf. [10, Theorem 4.4]. In particular, by [10, Definition 4.1], there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$. We get $\mathcal{B}_{r}(a)\cap\mathcal{H}\not=\emptyset$, which is also impossible. Thus, we conclude $\partial\mathcal{H}\cap F^{-1}(\{0\})=\{a_{+},a_{-}\}$. ∎
@@
All orbits in $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$ are unbounded. $\mathcal{H}$ is unbounded.
@@
If =$\partial\mathcal{H}=\{a_{+},a_{-}\}$, then $\mathcal{H}\not=\emptyset$ is open and closed with respect to the subspace topology in $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$. [^12] This argument is similar to the beginning of the proof of Proposition 4.5. Since $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is connected, =$\mathcal{H}=\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is clearly unbounded. Thus, assume $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}\not=\emptyset$.
Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. As in the proof of Theorem 2.6, $K:=\overline{\Gamma}$ is compact and $\Gamma\cup w_{\pm}(\Gamma)\subset K\subset\partial\mathcal{H}$. By Lemma 5.4, $\Gamma$ cannot be periodic. Hence, by Proposition 5.5 and [10, Theorem 5.1], we must have $w_{+}(\Gamma)\cup w_{-}(\Gamma)=\{a_{+},a_{-}\}$. Using this result, we want to derive a contradiction.
If $w_{+}(\Gamma)=\{a_{+}\}$ and $w_{-}(\Gamma)=\{a_{-}\}$, then we get $\Gamma\subset\mathcal{H}$, which is contraction to Theorem 5.3. If $w_{+}(\Gamma)=\{a_{-}\}$ and $w_{-}(\Gamma)=\{a_{+}\}$, then there exists another heteroclinic region $\tilde{\mathcal{H}}$ between $a_{-}$ and $a_{+}$. By Theorem 5.3, $\tilde{\mathcal{H}}$ is open, i.e. $\tilde{\mathcal{H}}\cap\mathcal{H}\not=\emptyset$, which is again a contradiction. Thus, $\Gamma$ must be homoclinic. Assume w.l.o.g. that $\Gamma$ is homoclinic in $a_{+}$, i.e. $a_{+}$ is an equilibrium of order $m\geq 2$. By Corollary 3.11, there exist two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma(x)$ is a sector-forming orbit in $a_{+}$ with these definite directions. By Theorem 3.6, we get $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))\cap\mathcal{H}\not=\emptyset$, i.e. there exists a homoclinic orbit in $\mathcal{H}$. This is a contradiction. We conclude that $\Gamma$ cannot be bounded. Moreover, since $\partial\mathcal{H}\not=\{a_{+},a_{-}\}$, there exists at least one unbounded orbit on $\partial\mathcal{H}$. Therefore, it follows that $\mathcal{H}$ is unbounded. ∎
@@
Suppose we find $p$ equilibria and $p$ heteroclinic orbits with these properties. Then $J:=a_{1}\cup\ldots\cup a_{p}\cup\Gamma_{1}\cup\ldots\cup\Gamma_{p}$ is a closed piecewise continuously differentiable Jordan curve. Let $\mathcal{H}_{j}$ be the heteroclinic region defined by $\Gamma_{j}$. Define the two open sets $U:=\mathcal{H}_{1}$ and $V:=\mathcal{H}_{2}\cup\ldots\cup\mathcal{H}_{p}$, cf. Theorem 5.3. Then there holds $U\cap V=\emptyset$ as well as, by Theorem 5.3, $U\cap\operatorname{Int}(J)\not=\emptyset$ and $V\cap\operatorname{Int}(J)\not=\emptyset$. Since $\operatorname{Int}(J)$ is connected, we get $U\cup V\not=\operatorname{Int}(J)$ and thus there exists $x\in\partial U\cap\operatorname{Int}(J)\not=\emptyset$. By Theorem 5.6, $\Gamma(x)\subset\partial\mathcal{H}_{1}\cap\operatorname{Int}(J)$ is unbounded. This is a contradiction. ∎
@@
$\mathcal{H}$ is connected, path-connected and simply connected.
@@
For $x\in\mathcal{H}$ define $A_{x}:=\Gamma(x)\cup\{a_{+}\}$ and the cover $C:=\{A_{x}:x\in\mathcal{H}\}$ of $\mathcal{H}$. Since $a_{+}$ lies on the boundary of each orbit in $\mathcal{H}$, every set in $C$ is connected, cf. [18, Theorem 23.4]. Moreover, for $A,\tilde{A}\in C$ it always holds $A\cap\tilde{A}\supset\{a_{+}\}\not=\emptyset$. Thus, by [18, Theorem 23.3], $\mathcal{H}$ is connected. By [23, Proposition 12.25], $\mathcal{H}$ is even path-connected.
Let $L\subset\mathcal{H}$ be a loop. Analogous to the beginning of the proof of Theorem 4.7, we find a closed Jordan curve $J\subset\mathcal{H}$ such that $L\subset\operatorname{Int}(J)$. Suppose that $\operatorname{Int}(J)\setminus\mathcal{H}\not=\emptyset$, i.e. $\partial\mathcal{H}\cap\operatorname{Int}(J)\not=\emptyset$.
Let $y\in\partial\mathcal{H}\cap\operatorname{Int}(J)$. By Proposition 5.2, we have $\Gamma(y)\subset\partial\mathcal{H}$. Since $\mathcal{H}$ is open, cf. Theorem 5.3, we get $\Gamma(y)\cap J=\emptyset$, i.e. $\Gamma(y)\subset\operatorname{Int}(J)$. By Proposition 5.5 and Theorem 5.6, $\Gamma(y)=\{y\}$ must be an equilibrium and thus $y\in\{a_{+},a_{-}\}$. If $y$ is an equilibrium of order $m\geq 2$, we can apply Corollary 3.11 to conclude the existence of an global elliptic sector in $y$ lying completely in $\operatorname{Int}(J)$, which is a contradiction to Theorem 3.7. Hence, by Lemma 5.4, $y$ must be a node or focus. If the second equilibrium $a\in\{a_{+},a_{-}\}\setminus\{y\}$ is also a node or focus, we can apply Proposition 4.5 and Theorem 4.6 to conclude the existence of an unbounded orbit $\Gamma_{0}\subset\partial\mathcal{N}$ with $y\in w_{+}(\Gamma_{0})\cup w_{-}(\Gamma_{0})$, where $\mathcal{N}$ is the basin in $a$, cf. Definition 4.1. Thus, $\Gamma_{0}\cap\mathcal{H}\supset\Gamma_{0}\cap J\not=\emptyset$, i.e. there exists an unbounded orbit in $\mathcal{H}$, which is impossible. It follows that $a$ must be an equilibrium of order $m\geq 2$ and therefore, as above, $a\in\operatorname{Ext}(J)$.
To summarise, if $\operatorname{Int}(J)\setminus\mathcal{H}\not=\emptyset$, the only possible case is that one of the two equilibria, $a_{+}$ and $a_{-}$, is a node or focus being the only point in $\operatorname{Int}(J)\setminus\mathcal{H}$ and the other is an equilibrium of order $m\geq 2$ in $\operatorname{Ext}(J)$. To fix ideas, assume w.l.o.g. $a_{+}\in\operatorname{Int}(J)$, $a_{-}\in\operatorname{Ext}(J)$ and $\operatorname{Int}(J)\setminus\mathcal{H}=\{a_{+}\}$. Using this and a similar construction as in the proof of Proposition 4.5, we want to derive a contradiction.
First, by [10, Theorem 4.4] and [10, Definition 4.1 c) (iii)], there exists a closed piecewise continuously differentiable Jordan curve $\Gamma$ that is nowhere tangential to $F$ and satisfies $\operatorname{Int}(\Gamma)\cap F^{-1}(\{0\})=\{a_{-}\}$. Moreover, all non-elliptic orbits in $a_{-}$, i.e. all orbits not lying in a global elliptic sector in $a_{-}$, cross $\Gamma$ exactly once. Hence, all orbits in $\mathcal{H}$ cross $C_{1}:=\Gamma\cap\mathcal{H}$ exactly once. We equip $C_{1}$ with the subspace topology. Using this construction, the map $\Psi:\mathcal{H}\to\mathbb{R}$ is well-defined, when $\Psi(x)$ is the transit time131313cf. [4, Definition 3.2]. the point $x\in\mathcal{H}$ needs to $C_{1}$ along the path $\Gamma(x)$, i.e. $\{\Phi(\Psi(x),x)\}=\Gamma(x)\cap C_{1}$.
Secondly, choose $\rho>0$ such that $\mathcal{B}_{\rho}(a_{+})\subset\operatorname{Int}(J)$. By applying the theory of circles without contact around nodes and foci, cf. [2, §3, 10.-14., §7, 1.-2. and §18, Lemma 3], we find a continuously differentiable closed path $C_{2}\subset\mathcal{B}_{\rho}(a)\subset\mathcal{H}$ that is nowhere tangential to $F$ and satisfies $\operatorname{Int}(C_{2})\cap F^{-1}(\{0\})=\{a_{+}\}$. Moreover, every orbit in $\mathcal{H}$ crosses $C_{2}$ exactly once, cf. [2, §3, 10., Figure 54]. We equip $C_{2}$ with the subspace topology and define the map $\eta:C_{2}\to C_{1}$ by $\eta(x):=\Phi(\Psi(x),x)$. Analogous to the proof of Proposition 4.5, $\eta$ is continuous and bijective. Additionally, $C_{1}$ is Hausdorff and $C_{2}$ is compact, i.e. $\eta$ is a homeomorphism, cf. [18, Theorem 26.6]. Furthermore, $C_{2}$ is isomorphic to $S^{1}$. By [18, Corollary 52.5] and [18, Theorem 54.5], it follows that the fundamental group of $C_{1}$ is isomorphic to the additive group $\mathbb{Z}$.
By [10, Definition 4.1 c) (ii)], there exists a homoclinic orbit having nonempty intersection with $\Gamma$. Thus, this homoclinic orbit cannot lie in $\mathcal{H}$. Therefore, it follows that $C_{1}$ cannot be closed, i.e. the fundamental group of $C_{1}$ cannot be isomorphic to $\mathbb{Z}$. More precisely, $C_{1}$ is connected and thus homeomorphic to the simply-connected space $(0,1)\subset\mathbb{R}$, cf. [18, Theorem 23.5]. This is a contradiction. All in all, we conclude $\operatorname{Int}(J)\subset\mathcal{H}$. Since $\operatorname{Int}(J)$ is simply connected, $L$ can be continuously deformed to a point. Since $L$ is arbitrary and $\mathcal{H}$ is already path-connected, it follows that $\mathcal{H}$ is even simply connected. ∎
In [3, 4, 10] many examples are illustrating topological properties of the basins of equilibria and global elliptic sectors. [10, Example 4.8, Figure 3] demonstrates a heteroclinic region between the two equilibria. The following example shows that there can be more than one heteroclinic region between two global elliptic sectors.

### v3

@@ 5. Heteroclinic regions
In this chapter we consider an entire function $F\in\mathcal{O}{}(\mathbb{C}{})$, $F\not\equiv 0$, with two equilibria $a_{+},a_{-}\in\mathbb{C}{}$ of (1.1) and at least one heteroclinic orbit connecting these two equilibria. By the holomorphy of $F$, the two equilibria have to be nodes, foci or multiple equilibria. Under these assumptions, it is possible to investigate topologically another characteristic set, the so-called "heteroclinic region". Taking up the questions at the end of chapter 3, a heteroclinic region is one possible structure between two adjacent global elliptic sectors. We verify that heteroclinic regions are strip canonical regions in the sense of Markus and Neumann, cf. [9, 11].
5.1. Geometry of heteroclinic regions
@@
$\mathcal{H}_{+(-)}:=\left\{x\in\mathbb{C}{}:w_{+(-)}(\Gamma(x))=\{a_{+(-)}\}\right\}\text{.}$
@@
Both, $a_{+}$ as well $a_{-}$ is either a focus or node or an equilibrium of order $m\in\mathbb{N}\setminus\{1\}$. Let $x\in\mathcal{H}_{+}$ and assume that $a_{+}$ is an equilibrium of order $m\geq 2$. By [19, Proposition 4.3], there exist numbers $r,\theta,\Theta,\tau>0$ such that for all $z\in\{\hat{z}\in\mathbb{C}{}:|\hat{z}-a_{+}|<r,|\arg(\hat{z}-a_{+})-\theta|<\Theta\}=:A$ the orbit through $z$ tends to $a$ for $t\to\infty$ and $M:=\Phi(\tau,x)\in A$. Since $A$ is open, we can choose $\varepsilon>0$ small enough such that $\mathcal{B}_{\varepsilon}(M)\subset A$. By continuity of the flow, cf. [29, Chapter 2.4, Theorem 4], there exists $\delta>0$ such that $|\Phi(\tau,x)-\Phi(\tau,y)|<\varepsilon$ for all $y\in\mathcal{B}_{\delta}(x)$, i.e. $\mathcal{B}_{\delta}(x)\subset\mathcal{H}_{+}$ and $\mathcal{H}_{+}$ is open. Additionally, if $a_{+}$ is a node or focus, then we can apply Theorem 4.3 to conclude again the openness of $\mathcal{H}_{+}$. Since the same argumentation holds for $a_{-}$, $\mathcal{H}_{-}$ is also open. It follows that $\mathcal{H}=\mathcal{H}_{+}\cap\mathcal{H}_{-}$ is open. ∎
@@
Since there exists at least one heteroclinic orbit in $\mathcal{H}$ connecting $a_{+}$ and $a_{-}$, we have $\{a_{+},a_{-}\}\subset\partial\mathcal{H}$. Suppose there exists an equilibrium $a\in\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. By Lemma 2.6, $a$ cannot be a center. Additionally, since all orbits in $\mathcal{H}$ connect $a_{+}$ to $_{-}, aa$ cannot be a node a focus. Hence, by using the holomorphy of $F$, we can apply [19, Theorem 4.4] to conclude that $a$ has a FED. In particular, there exists $r>0$ such that $a\in w_{+}(\Gamma(x))\cup w_{-}(\Gamma(x))$ for all $x\in\mathcal{B}_{r}(a)$. We get $\mathcal{B}_{r}(a)\cap\mathcal{H}\not$$\emptyset$, which is also impossible. Thus, we conclude $\partial\mathcal{H}\cap F^{-1}(\{0\})=\{a_{+},a_{-}\}$. ∎
@@
All orbits on $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$ are unbounded. $\mathcal{H}$ is unbounded.
@@
If $\partial\mathcal{H}=\{a_{+},a_{-}\}$, then $\mathcal{H}\not$$\emptyset$ is open and closed in the connected set $\mathbb{C}{}\setminus\{a_{+},a_{-}\}$, i.e. $\mathcal{H}=\mathbb{C}{}\setminus\{a_{+},a_{-}\}$ is clearly unbounded. Thus, we assume $\partial\mathcal{H}\setminus\{a_{+},a_{-}\}\not$$\emptyset$.
Suppose there exists a bounded orbit $\Gamma\subset\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. By Lemma 2.6, $\Gamma$ cannot be periodic. Hence, by Proposition 5.4 and Theorem 1.1, we must have $w_{+}(\Gamma)\cup w_{-}(\Gamma)\subset\{a_{+},a_{-}\}$. Additionally, by openness of $\mathcal{H}$, $\Gamma$ cannot be heteroclinic. It follows that $\Gamma$ must be homoclinic, w.l.o.g. in $a_{+}$, i.e. $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a_{+}\}$ and $a_{+}$ is an equilibrium of order $m\geq 2$. We choose a point $x\in\Gamma\subset\partial\mathcal{H}\setminus\{a_{+},a_{-}\}$. By Lemma 3.5, there exists $\delta>0$ such that all orbits $\hat{\Gamma}$ through $\mathcal{B}_{\delta}(x)$ are also homoclinic in $a_{+}$. Choosing $\hat{\Gamma}\in\mathcal{H}$ leads to a contradiction. We conclude that $\Gamma$ cannot be bounded. Moreover, since $\partial\mathcal{H}\not=\{a_{+},a_{-}\}$, there exists at least one unbounded orbit on $\partial\mathcal{H}$. Therefore, it follows that $\mathcal{H}$ is unbounded. ∎
@@
Suppose we find $p$ equilibria and $p$ heteroclinic orbits with these properties. Then $J:=a_{1}\cup\ldots\cup a_{p}\cup\Gamma_{1}\cup\ldots\cup\Gamma_{p}$ is a closed piecewise continuously differentiable Jordan curve. Let $\mathcal{H}_{j}$ be the heteroclinic region defined by $\Gamma_{j}$. Define the two open sets $U:=\mathcal{H}_{1}$ and $V:=\mathcal{H}_{2}\cup\ldots\cup\mathcal{H}_{p}$, cf. Theorem 5.3. Then there holds $U\cap V=\emptyset$ as well as, by Theorem 5.3, $U\cap\operatorname{Int}(J)\not$$\emptyset$ and $V\cap\operatorname{Int}(J)\not$$\emptyset$. Since $\operatorname{Int}(J)$ is connected, we get $U\cup V\not=\operatorname{Int}(J)$ and thus there exists $x\in\partial U\cap\operatorname{Int}(J)\not$$\emptyset$. By Theorem 5.5, $\Gamma(x)\subset\partial\mathcal{H}_{1}\cap\operatorname{Int}(J)$ is unbounded, which is a contradiction. ∎
This illustrates again that holomorphy severely constrains the possible geometric configurations of the phase space. In fact, Corollary 5.6 does not hold if $F$ is not holomorphic. This happens in particular when one of the equilibria in Corollary 5.6 is a saddle, cf. e.g. [13, Chapter 2].
Having established all of the above topological properties, we can now assemble them to reveal the full geometric structure of the heteroclinic region. The following argumentation shows the existence of a global transversal for a heteroclinic region.
  • (I)
    By using the local FED geometry in a multiple equilibrium, we deduce the existence of a transversal $T$ connecting two adjacent global elliptic sectors that have no common boundary near the multiple equilibrium, cf. [19, Definition 4.1] and Figure 1. Every orbit intersecting $T$, crosses it exactly once. Assume that a heteroclinic region $\mathcal{H}$ lies between these two sectors, and define the set $A=\mathcal{H}\cap T$. Let $q_{1},q_{2}\in A$, and consider the closed Jordan-curve $J$ formed by $\Gamma(q_{1})$, $\Gamma(q_{2})$, $a_{+}$ and $a_{-}$. Connecting $q_{1}$ and $a_{2}$ along the segment of $T$, one may ask whether this segment could leave $\mathcal{H}$? The answer is no. In fact, since the segment lies entirely in the interior of $J$, this would contradict Corollary 5.6 (with $p=2$). It follows that all points in $A$ are connected along the transversal $T$. Taking now the longest possible segment of the transversal $T$ contained in $\mathcal{H}$, we obtain an open smooth arc crossing each closed orbit in $\mathcal{H}$ exactly once. This is illustrated in Figure 3.
  • (II)
    Let $a$ be a focus or node on the boundary of a heteroclinic region $\mathcal{H}$. By applying the theory of circles without contact around nodes and foci, cf. [2, §3, 10.-14., §7, 1.-2. and §18, Lemma 3], we find always a continuously differentiable closed path $C\subset\mathcal{H}$ being nowhere tangential to $F$ and satisfying $\operatorname{Int}(C)\cap F^{-1}(\{0\})=\{a\}$. All orbits in the basin of $a$ cross $C$ exactly once. This is illustrated in Figure 4.
  • (III)
    If $a_{+}$ or $a_{-}$ is a multiple equilibrium, we can apply (I) to deduce the existence of global transversal. If both, $a_{+}$ and $a_{-}$ are foci/nodes, then we can apply Proposition 4.4 to show the existence of at least one unbounded orbit $\Gamma$ with $a_{+}$ in its limit set. Since this orbit intersects $C$ at exactly one point, cf. (II), $C\cap\mathcal{H}$ cannot be a closed segment of $C$, cf. Figure 4. Analogous to the argument in (I), $C\cap\mathcal{H}$ has to be path-connected and is thus a global transversal for $\mathcal{H}$.
Figure 3. A complete transversal (blue) for $\mathcal{H}$ near a multiple equilibrium (green). Separatrices are colored red.
Figure 4. A complete transversal (blue) for $\mathcal{H}$ near a node (green). The unbounded separatrix is colored red, cf. step (III).
@@
$\mathcal{H}$ is connected, path-connected and simply connected. $\mathcal{H}$ is a strip canonical region.
@@
We use our above considerations in (I)–(III). Let $T:[0,1]\to\mathcal{H}$ be the smooth global transversal through $\mathcal{H}$ and define $Q:=\mathbb{R}\times[0,1]\subset\mathbb{R}^{2}$. Then the map $\mu:Q\to\mathcal{H}$, $\mu(s,t):=\Phi(t,T(s))$ is a homeomorphism. Moreover, by openness of $\mathcal{H}$, there are no separatrices in $\mathcal{H}$. Since $Q$ is simply connected, this theorem is a direct consequence of [10, Theorem I] and [9, Theorem 5.1]. ∎
This leads to an interesting observation: global elliptic sectors and heteroclinic regions are both strip regions and thus topologically equivalent in the sense of [9]. In order to distinguish these two sets we need information about the boundary orbits (separatrices). This underlines that the behavior of separatrices is crucial for determining the global phase portrait of (1.1).
5.2. Illustrating example
Several examples of basins, sectors, and regions can be found in [4, 5, 19]. Another example has already been discussed, cf. Example 4.7. The vector field $F(x)=x^{3}-x$ has two heteroclinic regions: the left and right open half-planes of $\mathbb{C}{}\setminus\{x\in\mathbb{R}:|x|>1\}$. These heteroclinic regions share $\mathrm{i}\mathbb{R}$ (one equilibrium and two separatrices) as their common boundary. We now turn to a more complicated example to illustrate the theory developed in the preceding chapters.